OpenStudy (erinweeks):

Solve the system of equations using matrices. Use Gaussian elimination with back-substitution.

OpenStudy (erinweeks):

3x + 5y - 2w = -13 2x + 7z - w = -1 4y + 3z + 3w = 1 -x + 2y + 4z = -5

OpenStudy (erinweeks):

A. ( -1, 20/13, 0, 2/5 ) B. ( 1, -2, 0, 3 ) C. ( 3/4, -2, 0, 3/4 ) D. ( 4/3, - 13/20, 0, 5/2 )

ganeshie8 (ganeshie8):

can u write the equations in matrix form ?

OpenStudy (erinweeks):



OpenStudy (erinweeks):

i think?

ganeshie8 (ganeshie8):

careful, u have 4 variables, so u must have 4 clomns aswell

ganeshie8 (ganeshie8):

\[ \left[ \begin{array}{ccc} 3 & 5 & 0 & -2 \\ 2 & 0 & 7 & -1 \\ 0 & 4 & 3 & 3 \\ -1 & 2 & 4 & 0 \\ \end{array} \right] \left[ \begin{array}{ccc} -13 \\ -1 \\ 1 \\ -5 \\ \end{array} \right] \]

ganeshie8 (ganeshie8):

reduce it to triangular matrix

OpenStudy (erinweeks):

oh sorry ! im not good at this still lol

OpenStudy (erinweeks):

i dont know how to reduce?

ganeshie8 (ganeshie8):

R1-R2

ganeshie8 (ganeshie8):

subtract R2 from R1

ganeshie8 (ganeshie8):

\[ \left[ \begin{array}{ccc} 3 & 5 & 0 & -2 \\ 2 & 0 & 7 & -1 \\ 0 & 4 & 3 & 3 \\ -1 & 2 & 4 & 0 \\ \end{array} \right] \left[ \begin{array}{ccc} -13 \\ -1 \\ 1 \\ -5 \\ \end{array} \right] \] R1-R2 \[ \left[ \begin{array}{ccc} 1 & 5 & -7 & -1 \\ 2 & 0 & 7 & -1 \\ 0 & 4 & 3 & 3 \\ -1 & 2 & 4 & 0 \\ \end{array} \right] \left[ \begin{array}{ccc} -12 \\ -1 \\ 1 \\ -5 \\ \end{array} \right] \]

ganeshie8 (ganeshie8):

ive just subtracted R2 from R1

ganeshie8 (ganeshie8):

see if u okays wid that

OpenStudy (erinweeks):

okay now what though?

ganeshie8 (ganeshie8):

next, we want to make first column as below :- 1 0 0 0

ganeshie8 (ganeshie8):

for that we do this :- R2-2R1 R4+R1

OpenStudy (erinweeks):

now im confused can you show me lol!

ganeshie8 (ganeshie8):

\[ \left[ \begin{array}{ccc} 3 & 5 & 0 & -2 \\ 2 & 0 & 7 & -1 \\ 0 & 4 & 3 & 3 \\ -1 & 2 & 4 & 0 \\ \end{array} \right] \left[ \begin{array}{ccc} -13 \\ -1 \\ 1 \\ -5 \\ \end{array} \right] \] R1-R2 \[ \left[ \begin{array}{ccc} 1 & 5 & -7 & -1 \\ 2 & 0 & 7 & -1 \\ 0 & 4 & 3 & 3 \\ -1 & 2 & 4 & 0 \\ \end{array} \right] \left[ \begin{array}{ccc} -12 \\ -1 \\ 1 \\ -5 \\ \end{array} \right] \] R2-2R1 R4+R1 \[ \left[ \begin{array}{ccc} 1 & 5 & -7 & -1 \\ 0 & -10 & 21 & 0 \\ 0 & 4 & 3 & 3 \\ 0 & 7 & -3 & -1 \\ \end{array} \right] \left[ \begin{array}{ccc} -12 \\ 23 \\ 1 \\ -17 \\ \end{array} \right] \]

OpenStudy (erinweeks):

okay i get it

ganeshie8 (ganeshie8):

\[ \left[ \begin{array}{ccc} 3 & 5 & 0 & -2 \\ 2 & 0 & 7 & -1 \\ 0 & 4 & 3 & 3 \\ -1 & 2 & 4 & 0 \\ \end{array} \right] \left[ \begin{array}{ccc} -13 \\ -1 \\ 1 \\ -5 \\ \end{array} \right] \] R1-R2 \[ \left[ \begin{array}{ccc} 1 & 5 & -7 & -1 \\ 2 & 0 & 7 & -1 \\ 0 & 4 & 3 & 3 \\ -1 & 2 & 4 & 0 \\ \end{array} \right] \left[ \begin{array}{ccc} -12 \\ -1 \\ 1 \\ -5 \\ \end{array} \right] \] R2-2R1 R4+R1 \[ \left[ \begin{array}{ccc} 1 & 5 & -7 & -1 \\ 0 & -10 & 21 & 0 \\ 0 & 4 & 3 & 3 \\ 0 & 7 & -3 & -1 \\ \end{array} \right] \left[ \begin{array}{ccc} -12 \\ 23 \\ 1 \\ -17 \\ \end{array} \right] \] R2/-10 \[ \left[ \begin{array}{ccc} 1 & 5 & -7 & -1 \\ 0 & 1 & -2.1 & 0 \\ 0 & 4 & 3 & 3 \\ 0 & 7 & -3 & -1 \\ \end{array} \right] \left[ \begin{array}{ccc} -12 \\ -2.3 \\ 1 \\ -17 \\ \end{array} \right] \]

ganeshie8 (ganeshie8):

next, ive divided R2 by -10

OpenStudy (erinweeks):

why did you do divide by 10?

ganeshie8 (ganeshie8):

cuz i want 1 at 22 position (row2, col2)

OpenStudy (erinweeks):

okay gotcha!! just was confused for a sec

ganeshie8 (ganeshie8):

\[ \left[ \begin{array}{ccc} 3 & 5 & 0 & -2 \\ 2 & 0 & 7 & -1 \\ 0 & 4 & 3 & 3 \\ -1 & 2 & 4 & 0 \\ \end{array} \right] \left[ \begin{array}{ccc} -13 \\ -1 \\ 1 \\ -5 \\ \end{array} \right] \] R1-R2 \[ \left[ \begin{array}{ccc} 1 & 5 & -7 & -1 \\ 2 & 0 & 7 & -1 \\ 0 & 4 & 3 & 3 \\ -1 & 2 & 4 & 0 \\ \end{array} \right] \left[ \begin{array}{ccc} -12 \\ -1 \\ 1 \\ -5 \\ \end{array} \right] \] R2-2R1 R4+R1 \[ \left[ \begin{array}{ccc} 1 & 5 & -7 & -1 \\ 0 & -10 & 21 & 0 \\ 0 & 4 & 3 & 3 \\ 0 & 7 & -3 & -1 \\ \end{array} \right] \left[ \begin{array}{ccc} -12 \\ 23 \\ 1 \\ -17 \\ \end{array} \right] \] R2/-10 \[ \left[ \begin{array}{ccc} 1 & 5 & -7 & -1 \\ 0 & 1 & -2.1 & 0 \\ 0 & 4 & 3 & 3 \\ 0 & 7 & -3 & -1 \\ \end{array} \right] \left[ \begin{array}{ccc} -12 \\ -2.3 \\ 1 \\ -17 \\ \end{array} \right] \] R3-4R2 R4-7R2 \[ \left[ \begin{array}{ccc} 1 & 5 & -7 & -1 \\ 0 & 1 & -2.1 & 0 \\ 0 & 0 & 11.4 & 3 \\ 0 & 0 & 11.7 & -1 \\ \end{array} \right] \left[ \begin{array}{ccc} -12 \\ -2.3 \\ 10.2 \\ -0.9 \\ \end{array} \right] \]

ganeshie8 (ganeshie8):

plz do check if my calculations are correct or not :)

OpenStudy (erinweeks):

i believe they are!!! so is that it..

OpenStudy (erinweeks):

or what do we do from here?

ganeshie8 (ganeshie8):

no, we need to conver it upper triangulear

OpenStudy (erinweeks):

i didnt really get that.. that doesnt look like any of my answers

ganeshie8 (ganeshie8):

last column is the solution

OpenStudy (erinweeks):

that doesnt look like this though A. ( -1, 20/13, 0, 2/5 ) B. ( 1, -2, 0, 3 ) C. ( 3/4, -2, 0, 3/4 ) D. ( 4/3, - 13/20, 0, 5/2 )

OpenStudy (erinweeks):

oh okay so its always the last column ?

ganeshie8 (ganeshie8):

yes